right-tailed test about population variance o2, where n = 18 and a = 0.01 (10)
Answer #1

solution Given that n =18 degrees of Freedom = n.1=18-1=17 x = 0.01 I -2,df xx²=1-0.01 17 (right tailed) x = 0.99.17 = 6.408